OpenStudy (anonymous):

A rectangular park, with dimensions of 1500ft x 2000ft has a diagonal walking path that goes from corner to corner. How long is the walking path? 250ft 2500ft 3500ft 5000ft @Directrix

OpenStudy (compassionate):

\[Diagonal = \sqrt{w^2 + h^2}\]\[Diagonal = \sqrt{1500^2 + 2000^2}\]

OpenStudy (anonymous):

wrong pic

OpenStudy (anonymous):

@Compassionate i got the answer of 6,250,000........ @Directrix can u help?

Directrix (directrix):